# G-VEval: A Versatile Metric for Evaluating Image and Video Captions Using GPT-4o

### v1

@@ Abstract
Evaluation metric of visual captioning is important yet not thoroughly explored. Traditional metrics like BLEU, METEOR, CIDEr, and ROUGE often miss semantic depth, while trained metrics such as CLIP-Score, PAC-S, and Polos are limited in zero-shot scenarios. Advanced Language Model-based metrics also struggle with aligning to nuanced human preferences. To address these issues, we introduce G-VEval, a novel metric inspired by G-Eval and powered by the new GPT-4o. G-VEval uses chain-of-thought reasoning in large multimodal models and supports three modes: reference-free, reference-only, and combined, accommodating both video and image inputs. We also propose MSVD-Eval, a new dataset for video captioning evaluation, to establish a more transparent and consistent framework for both human experts and evaluation metrics. It is designed to address the lack of clear criteria in existing datasets by introducing distinct dimensions of Accuracy, Completeness, Conciseness, and Relevance (ACCR). Extensive results show that G-VEval outperforms existing methods in correlation with human annotations, as measured by Kendall tau-b and Kendall tau-c. This provides a flexible solution for diverse captioning tasks and suggests a straightforward yet effective approach for large language models to understand video content, paving the way for advancements in automated captioning.  [^1] https://github.com/ztangaj/gveval
@@ 5 Discussion
However, G-VEval has certain limitations. One major drawback is its cost. While GPT-4o is relatively more affordable than some alternatives, it remains more expensive than other metrics due to the token-based pricing model.
Another potential concern is the consistency of scoring over time. Although current results demonstrate consistent performance, future updates to the GPT-4o model or changes in prompts could affect this consistency.
@@ 6 Conclusion
Looking ahead, we aim to develop an online benchmark platform based on G-VEval, where researchers can evaluate their image and video captioning models. This platform will pave the way for future research in image and video captioning tasks, providing a valuable resource for advancing the field and improving the quality of automated captioning systems.
@@ Appendix C Appendix C: MSVD-Eval Dataset
The following are some samples from the MSVD-Eval dataset, as shown in Figure 4:
@@
Appendix D Appendix D: Sample Qualitative Outputs
Here we provide sample qualitative outputs from Flickr-8k, VATEX-Eval, and MSVD-Eval.
D.1 S
ee Figure 5
@@ D.2 VATEX-Eval
See Figure 6
@@ D.3 MSVD-Eval
Figure 7: Samples from the MSVD-Eval dataset

### v2

@@ Abstract
Evaluation metric of visual captioning is important yet not thoroughly explored. Traditional metrics like BLEU, METEOR, CIDEr, and ROUGE often miss semantic depth, while trained metrics such as CLIP-Score, PAC-S, and Polos are limited in zero-shot scenarios. Advanced Language Model-based metrics also struggle with aligning to nuanced human preferences. To address these issues, we introduce G-VEval, a novel metric inspired by G-Eval and powered by the new GPT-4o. G-VEval uses chain-of-thought reasoning in large multimodal models and supports three modes: reference-free, reference-only, and combined, accommodating both video and image inputs. We also propose MSVD-Eval, a new dataset for video captioning evaluation, to establish a more transparent and consistent framework for both human experts and evaluation metrics. It is designed to address the lack of clear criteria in existing datasets by introducing distinct dimensions of Accuracy, Completeness, Conciseness, and Relevance (ACCR). Extensive results show that G-VEval outperforms existing methods in correlation with human annotations, as measured by Kendall tau-b and Kendall tau-c. This provides a flexible solution for diverse captioning tasks and suggests a straightforward yet effective approach for large language models to understand video content, paving the way for advancements in automated captioning.
Codehttps://github.com/ztangaj/gveval
Published versionTBD
@@ 2.3 Advanced Language Model-Based Metrics
Most recently, FLEUR proposes a novel evaluation framework by utilizing LLaVA, a vision-language model that can directly assess image-caption pairs (Lee, Park, and Kang 2024). Unlike CLAIR’s reference-dependent approach, FLEUR enables direct comparison between images and captions. It introduces a score smoothing technique that utilizes probability distributions for fine-grained evaluation. To improve computational efficiency, FLEUR separates score generation from explanation production, where explanations serve as post-hoc justifications through additional prompts after scoring.
@@ 5 Discussion
However, G-VEval has certain limitations. One major drawback is its cost. While GPT-4o is relatively more affordable than some alternatives, it remains more expensive than other metrics due to the token-based pricing model. Another potential concern is the consistency of scoring over time. Although current results demonstrate consistent performance, future updates to the GPT-4o model or changes in prompts could affect this consistency.
@@ 6 Conclusion
Looking ahead, we aim to develop an online benchmark platform based on G-VEval, where researchers can evaluate their image and video captioning models, furthering research and innovation in automated visual understanding.
@@ Appendix C Appendix C: MSVD-Eval Dataset
The following are some samples from the MSVD-Eval dataset, as shown in Figure 4, evaluated under our ACCR framework where Acc. , Com. , Con. , and Rel. represent Accuracy, Completeness, Conciseness, and Relevance scores respectively.
@@
Appendix D Appendix D: Sample Qualitative Outputs of G-VEval
Here we provide sample G-VEval evaluation outputs from Flickr-8k, VATEX-EVAL, and MSVD-Eval.
D.1 Flickr-8k
See Figure 5 for G-VEval’s evaluation of image captions.
@@ D.2 VATEX-EVAL
See Figure 6 for G-VEval’s evaluation of video captions.
@@ D.3 MSVD-Eval
See Figure 7 and 8 for G-VEval’s evaluation results under the ACCR framework.
Figure 7: Sample qualitative output from the MSVD-Eval dataset (Part 1)
Figure 8: Sample qualitative output from the MSVD-Eval dataset (Part 2)